OpenStudy (anonymous):

Please tell me how to rationalize the denominator for this problem 4x/√7x-2

OpenStudy (anonymous):

is it like this? \[\frac{ 4x }{ \sqrt{7x-2} }\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

then multiply top and bottom by the denominator...\[\frac{ 4x }{ \sqrt{7x-2} }=\frac{ 4x }{ \sqrt{7x-2} }\cdot \frac{ \sqrt{7x-2} }{ \sqrt{7x-2} }\]

OpenStudy (anonymous):

so long as \[x \ne \frac{ 2 }{ 7 }\]

OpenStudy (anonymous):

what would the answer to 6√125 x^15y^3 be? please this is my last question and im so frustrated

OpenStudy (anonymous):

? what is the question?

OpenStudy (anonymous):

\[\sqrt[6]{125x^{15}y ^{3}}\]

OpenStudy (anonymous):

are you supposed to simplify?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt[6]{125x^{15}y^3}=\sqrt[6]{5^3x^{15}y^3}=5^{\frac{ 3 }{ 6 }}x^{\frac{ 15 }{ 6 }}y^{\frac{ 3 }{ 6 }} = 5^{\frac{ 1 }{ 2 }}x^{2\frac{ 1 }{ 2 }}y^{\frac{ 1 }{ 2 }}=x^2\sqrt{5xy}\]

OpenStudy (anonymous):

Thank you sooo much! Seriously

OpenStudy (anonymous):

you're welcome, seriously